A 300 g glass thermometer initially at 21 ◦C
is put into 247 cm3
of hot water at 81 ◦C.
Find the final temperature of the thermometer, assuming no heat flows to the
surroundings. The specific heat of glass is
0.2 cal/g · ◦ C and of water 1 cal/g · ◦ C.
1
Expert's answer
2020-04-02T11:10:16-0400

The final temperature in the system after mixing

"t=\frac{c_1m_1t_1+c_2m_2t_2}{c_1m_1+c_2m_2}."

"=\frac{0.2\times 300\times 21+1.0\times 247\times 81}{0.2\times 300+1.0\times 247}=69\:^{\circ}\rm C."